OpenStudy (anonymous):

Graph the system of equations x + y = 7 2x - 4y = 8 My graphing calculator doesn't make any sense, how do I solve this without it?

myininaya (myininaya):

Solve both of those equations for y. Objective: Put into y=mx+b form.

OpenStudy (anonymous):

y = -x + 7 y = -.5x - 2

myininaya (myininaya):

I'm not sure about your second equation. I think you should have a positive slope.

myininaya (myininaya):

y=-x+7 y=x/2-2 Now graph using the y-intercepts and the slopes of each equation. y=mx+b (0,b) is the y-intercept m is the slope

OpenStudy (anonymous):

what exactly am i even supposed to be finding though? the x and y intercepts of each line?

myininaya (myininaya):

Well you said you wanted to solve the system of equations, right? So you are looking for when the lines meet. At what point, do they meet?

OpenStudy (anonymous):

(6, 1)

myininaya (myininaya):

Yep. That works. You can always check yourself by pluggin' in your result into both equations to see if both equations hold. Like so: (x=6,y=1) y=-x+7 1=-6+7 1=1 (1=1 is a true statement so (6,1) works for this equation but what about the other) y=x/2-2 1=6/2-2 1=3-2 1=1 (again this is a true statement) Since both of our equations hold, that means we have found the correct solution.

OpenStudy (anonymous):

i have another one that i'm trying to figure out too. this one says to solve it algebraically (with substitution or elimination) 5x + y = 9 6x + 10y = 2

OpenStudy (anonymous):

if i do it with substitution i get this 5x + y = 9 solved for y is y = -5x + 9 6x + 10(-5x + 9) = 2 6x - 50x + 9 = 2 6x - 50x = -7 -44x = -7 x = .159 which doesnt seem right at all, but if i try it with elimination then i get this 5x + y = 9 6x + 10y = 2 -10(5x + y = 9) = -50x - 10y = -90 -50x - 10y = -90 + 6x + 10y = 2 ---------------------- -44x = -88 x = 2 which seems right so, if x = 2, then the equations now look like this 10 + y = 9, and 12 + 10y = 2 y = -1 and 10y = -10, or y = -1

myininaya (myininaya):

well 10(9)=90